Substitution:
14. y=-3x+4
-6x-2y=-8
16. x=y-1
-x+y=-1
18. y=-3x+1
2x+y=-4
20. 5x-y=5
-x+3y=13
22. -5x+4y=20
10x-8y=-40
Please Help! Help Needed ASAP! PLEASE

Answer by Alan3354(69443)

Substitution:
14. y=-3x+4
-6x-2y=-8
16. x=y-1
-x+y=-1
18. y=-3x+1
2x+y=-4
20. 5x-y=5
-x+3y=13
22. -5x+4y=20
10x-8y=-40
===========================
Is that all 1 problem?